OpenStudy (anonymous):

functions f an g defined as f(x)=2x-5 an g(x)=x[2]+3 what is f(4) gf(4) an the inverse of f (x)

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

is this the function for g? \[g(x) = x^{2}+3\]

OpenStudy (anonymous):

i dont know that why am askin

OpenStudy (anonymous):

Anjalie: What joemath is doing is asking if that's what you meant when you posted the question. What you wrote actually doesn't make mathematical sense... You'll have to help him understand what the question looks like if you want him to help you.

OpenStudy (anonymous):

or okay yes thank you

OpenStudy (anonymous):

im just gonna solve it with what im assuming the function to be. First problem: To compute f(4), you take the function f, and everywhere you see an x, plug in a 4: \[f(x) = 2x-5 \Rightarrow f(4) = 2(4)-5 \Rightarrow f(4) = 8-5 = 3 \] So f(4) = 3.

OpenStudy (anonymous):

Gonna assume the gf(4) means g(f(4)) (although it could mean something else). We already figured out that f(4) = 3, so we just need g(3): \[g(3) = (3)^{2}+3 = 9+3 = 12\] so gf(4) = 12

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

For the inverse of f, change the 'f(x)' to a 'y', and solve for x: \[f(x) = 2x-5 \Rightarrow y = 2x-5 \Rightarrow y+5 = 2x\Rightarrow \frac{y+5}{2} = x\] Thus f inverse is: \[f^{-1}(y) = \frac{y+5}{2}\]

OpenStudy (anonymous):

ok thank you so much